OpenStudy (anonymous):

What is the reciprocal of –5?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

-1/5

OpenStudy (anonymous):

I think it would be -1/5

OpenStudy (anonymous):

i got it first

OpenStudy (anonymous):

oohooh pick me!

OpenStudy (some.random.cool.kid):

yes great job you are both correct :}

OpenStudy (anonymous):

Explain your answer...

OpenStudy (some.random.cool.kid):

sat ill give you medal xD

OpenStudy (anonymous):

under every number is a invisible 1 for fractions and whole number math

OpenStudy (anonymous):

in a multiplication or division question the whole number would be reciprocoled makeing it 1 over the whole number

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

no probs

OpenStudy (anonymous):

remember any reciprocol is just the whole number under 1

OpenStudy (anonymous):

uness its a mixed fraction

OpenStudy (anonymous):

Yeah

OpenStudy (solomonzelman):

The reciprocal of \(\large\color{royalblue}{ \rm \frac{1}{3} }\) ? \(\large\color{royalblue}{ \rm \frac{1}{3}\times b=1 }\) (multiply both sides times 3, and you get) \(\large\color{royalblue}{ \rm b=3 }\) so the reciprocal of \(\large\color{royalblue}{ \rm \frac{1}{3} }\), is \(\large\color{royalblue}{ \rm 3 }\) . ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ The reciprocal of \(\large\color{royalblue}{ \rm -\frac{1}{8} }\) ? \(\large\color{royalblue}{ \rm -\frac{1}{8}\times b=1 }\) (multiply both sides times -8, and you get) \(\large\color{royalblue}{ \rm b=-8 }\) so the reciprocal of \(\large\color{royalblue}{ \rm \frac{1}{3} }\), is \(\large\color{royalblue}{ \rm 3 }\) . ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ The reciprocal of \(\large\color{royalblue}{ \rm 4 }\) ? \(\large\color{royalblue}{ \rm 4\times b=1 }\) (divide both sides by 4, and you get) \(\large\color{royalblue}{ \rm \frac{1}{4}}\) so the reciprocal of \(\large\color{royalblue}{ \rm 4 }\), is \(\large\color{royalblue}{ \rm \frac{1}{4}}\) . ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ The reciprocal of \(\large\color{royalblue}{ \rm -5 }\) ? \(\large\color{royalblue}{ \rm -5\times b=1 }\) (divide both sides by -5, and you get) \(\large\color{royalblue}{ \rm -\frac{1}{5}}\) so the reciprocal of \(\large\color{royalblue}{ \rm -5 }\), is \(\large\color{royalblue}{ \rm -\frac{1}{5}}\) . ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

OpenStudy (solomonzelman):

These are just some examples.